# compat=1.1.0

### v1

@@ IV.2 Dark Radiation Constraints
×$\times\Delta N_{e}^{\rho}\left[\frac{T(N_{e}^{\rho})}{H(N_{e}^{\rho})}\right]\left[\frac{T(N_{e}^{\rho})}{M_{\rm pl}}\right]^{2}.$ (27)
While the precise value of $\Delta N_{\rm eff}$ depends on the detailed evolution between the peak in GW energy density production, at $N_{e}^{\rho}$, and the onset of radiation domination, at $N_{e}^{\rm RD}$, we can estimate a conservative upper bound by setting $[T(N_{e}^{\rho})/T(N_{e}^{\rm RD})\cdot e^{N_{e}^{\rho}-N_{e}^{\rm RD}}]=1$. This approximation assumes a smooth transition to the radiation-dominated phase after the end of inflation, where both the bath temperature and the Hubble parameter are monotonically decreasing functions of e-folds. In these simple scenarios, the temperature evolves more gradually than in radiation domination due to the ongoing entropy injection from the inflaton decay, which slows the cooling relative to standard adiabatic expansion.
@@
×$\times\frac{1}{n^{\frac{3}{2}}\left(1+4\,N_{e}^{\rm cmb}/n\right)^{\frac{n+2}{8}}},$ (28)
@@ 1 An upper bound on the contribution of thermal GWs to the primordial tensor spectrum
[…]
@@
×$\times\left[\frac{n^{2}}{1+Q}\,\left(1+\frac{4\Delta N_{e}^{\rm ex}}{n}\right)\right]^{\frac{n-2}{8}}\,M_{\rm pl}.$ (35)
@@
×$\times\,\frac{1}{n\left(n+4\,\Delta N_{e}^{\rm ex}\right)}\frac{Q}{1+Q}.$ (37)
@@ 2 An upper bound on the contribution of thermal GWs to dark radiation
We now evaluate the contribution to $\Delta N_{\rm eff}$ from thermal GWs, Eq (27). Assuming a smooth transition to the radiation-dominated phase after the end of inflation, where both the bath temperature $T$ and Hubble parameter $H$ are monotonically decreasing functions of e-folds, we can estimate a conservative upper bound on $\Delta N_{\rm eff}$ by: (i) setting $[T(N_{e}^{\rho})/T(N_{e}^{\rm RD})\cdot e^{N_{e}^{\rho}-N_{e}^{\rm RD}}]=1$ and (ii) evaluating it based on the values of $T$ and $H$ at the end of inflation rather than at the peak of the GW production $N_{e}^{\rm\rho}$, since both $T$ and $T/H$ will be greater at this earlier time. In formulas, we can write:
@@
×$\times\frac{Q^{\frac{3}{4}}}{(1+Q)^{\frac{n+6}{8}}}\,\lambda_{n}^{\frac{1}{4}}.$ (41)
@@
×$\times\left(\frac{1}{1+4N_{e}^{\rm cmb}/n}\right)^{\frac{n+2}{8}},$ (45)
@@
×$\times\frac{1}{n^{\frac{3}{2}}\left(1+4\,N_{e}^{\rm cmb}/n\right)^{\frac{n+2}{8}}}.$ (46)

### v2

@@ IV.2 Dark Radiation Constraints
$\times\Delta N_{e}^{\rho}\left[\frac{T(N_{e}^{\rho})}{H(N_{e}^{\rho})}\right]\left[\frac{T(N_{e}^{\rho})}{M_{\rm pl}}\right]^{2}.$ (27)
While the precise value of $\Delta N_{\rm eff}$ depends on the detailed evolution between the peak in GW energy density production, at $N_{e}^{\rho}$, and the onset of radiation domination, at $N_{e}^{\rm RD}$, we can estimate a conservative upper bound by setting $T(N_{e}^{\rho})/T(N_{e}^{\rm RD})\cdot e^{N_{e}^{\rho}-N_{e}^{\rm RD}}]=1$. This approximation assumes a smooth transition to the radiation-dominated phase after the end of inflation, where both the bath temperature and the Hubble parameter are monotonically decreasing functions of e-folds. In these simple scenarios, the temperature evolves more gradually than in radiation domination due to the ongoing entropy injection from the inflaton decay, which slows the cooling relative to standard adiabatic expansion.
@@
$\times\frac{1}{n^{\frac{3}{2}}\left(1+4\,N_{e}^{\rm cmb}/n\right)^{\frac{n+2}{8}}},$ (28)
@@ 1 An upper bound on the contribution of thermal GWs to the primordial tensor spectrum
$\Delta N_{e}^{\rm ex}=\frac{1+Q}{M_{\rm pl}}\int_{\tilde{\phi}(N_{e}^{\rm ex})}^{\tilde{\phi}_{\rm end}}\frac{V_{,\phi}}{V}d\tilde{\phi}$, (33)
@@
$\times\left[\frac{n^{2}}{1+Q}\,\left(1+\frac{4\Delta N_{e}^{\rm ex}}{n}\right)\right]^{\frac{n-2}{8}}\,M_{\rm pl}.$ (35)
@@
$\times\,\frac{1}{n\left(n+4\,\Delta N_{e}^{\rm ex}\right)}\frac{Q}{1+Q}.$ (37)
@@ 2 An upper bound on the contribution of thermal GWs to dark radiation
We now evaluate the contribution to $\Delta N_{\rm eff}$ from thermal GWs, Eq (27). Assuming a smooth transition to the radiation-dominated phase after the end of inflation, where both the bath temperature $T$ and Hubble parameter $H$ are monotonically decreasing functions of e-folds, we can estimate a conservative upper bound on $\Delta N_{\rm eff}$ by: (i) setting $T(N_{e}^{\rho})/T(N_{e}^{\rm RD})\cdot e^{N_{e}^{\rho}-N_{e}^{\rm RD}}]=1$ and (ii) evaluating it based on the values of $T$ and $H$ at the end of inflation rather than at the peak of the GW production $N_{e}^{\rm\rho}$, since both $T$ and $T/H$ will be greater at this earlier time. In formulas, we can write:
@@
$\times\frac{Q^{\frac{3}{4}}}{(1+Q)^{\frac{n+6}{8}}}\,\lambda_{n}^{\frac{1}{4}}.$ (41)
@@
$\times\left(\frac{1}{1+4N_{e}^{\rm cmb}/n}\right)^{\frac{n+2}{8}},$ (45)
@@
$\times\frac{1}{n^{\frac{3}{2}}\left(1+4\,N_{e}^{\rm cmb}/n\right)^{\frac{n+2}{8}}}.$ (46)